find 3 consecutive integers so that twice the first is twelve more than the third.
Answer by Kalmetam(43)

so we are finding:
A,B,C
and those are the numbers from least to greatest..
and:
2B=N
C+12=N
Lets start RANDOM!!
lets say:
1,2,3
Then
2*2=4
4=n
C+12=n FALSE
---
Lets try a bit higher..
---
11,12,13
12*2=24
24=n
13+12=24 WRONG
----
That was extremely close
lets try one number up and exclude the 11.. that should work since 24 is one number down..
----
12,13,14
13*2=26
14+12=26 TRUE
---
The integers are: 12,13,14